### v3

# ACEBench: Who Wins the Match Point in Tool Usage?

Abstract
Large Language Models (LLMs) have demonstrated significant potential in decision-making and reasoning, particularly when integrated with various tools to effectively solve complex problems. However, existing benchmarks for evaluating LLMs’ tool usage face several limitations: (1) limited evaluation scenarios, often lacking assessments in real multi-turn dialogue contexts; (2) narrow evaluation dimensions, with insufficient detailed assessments of how LLMs use tools; and (3) reliance on LLMs or real API executions for evaluation, which introduces significant overhead. To address these challenges, we introduce ACEBench, a comprehensive benchmark for assessing tool usage in LLMs. ACEBench categorizes data into three primary types based on evaluation methodology: Normal, Special, and Agent. "Normal" evaluates tool usage in basic scenarios; "Special" evaluates tool usage in situations with ambiguous or incomplete instructions; "Agent" evaluates tool usage through multi-agent interactions to simulate real-world, multi-turn dialogues. We conducted extensive experiments using ACEBench, analyzing various LLMs in-depth and providing a more granular examination of error causes across different data types.
ACEBench: Who Wins the Match Point in Tool Usage?
11footnotetext: Equal Contributions. Work was done during Chen Chen’s internship at Huawei Noah’s Ark Lab22footnotetext: Corresponding authors.
1 Introduction
Large Language Models (LLMs), such as GPT-4 Achiam et al. (2023), have demonstrated exceptional performance across numerous natural language processing tasks Naveed et al. (2023); Qu et al. (2025); Mialon et al. (2023). Studies have shown that incorporating tools can significantly expand LLM capabilities, particularly in specialized domains such as mathematics Das et al. (2024); Bulusu et al. (2024); Gou et al. (2023); Veerendranath et al. (2024), programming Xu et al. (2024), and reasoning Chen et al. (2022); Shao et al. (2022); Surís et al. (2023); Yang et al. (2023). On one hand, integrating tools into LLMs can enhance capabilities in multiple domains, for example, ToolTransformer Schick et al. (2023) enhances the ability of LLMs to solve complex problems by utilizing tools. On the other hand, adopting a tool usage paradigm can improve response robustness and the transparency of the generation, thereby increasing explainability and user trust Schick et al. (2023), as well as improving the system’s adaptability. As this field continues to evolve, it is essential to comprehensively evaluate all aspects of tool usage, particularly in complex scenarios.
Table 1: Comparison of benchmarks across different evaluation criteria. "LLM-Free" refers to result evaluation without relying on LLMs. "Robustness" refers to incomplete or unclear user instructions. "Interactiveness" refers to the dynamic interaction between the model and the environment. "Atomic-Level" refers to analyzing from the atomic-level capabilities. "Personalization” refers to the inclusion of personal likes.
Benchmark LLM-Free Robustness Interactiveness Atomic-Level Personalization
MetaTool Huang et al. (2023)
API-Bank Li et al. (2023)
Stable ToolBench Guo et al. (2024)
BFCL Yan et al. (2024)
TauBench Alanko et al. (2022)
HammerBench Wang et al. (2024a)
ACEBench (Ours)
While several studies have focused on evaluating tool usage Yan et al. (2024); Guo et al. (2024); Wang et al. (2024a); Qin et al. (2023); Wang et al. (2024b); Zhuang et al. (2023); Lu et al. (2024), there are still some shortcomings in the existing tool-use benchmarks. Firstly, existing benchmarks lack multi-turn dialogue evaluation in real-world scenarios. For example, the multi-turn dialogues in BFCL Yan et al. (2024) and HammerBench  Wang et al. (2024a) are composed of predefined fixed content combinations. Secondly, current tool-use benchmarks Qin et al. (2023); Guo et al. (2024); Huang et al. (2023); Li et al. (2023) lack fine-grained evaluation and personalized data assessment. Additionally, existing benchmarks Qin et al. (2023); Guo et al. (2024); Wang et al. (2024b) ignore the assessment of special cases, or the evaluation methods are simplistic Yan et al. (2024), as user instructions in real life are not always perfectWang et al. (2024c). The model’s ability to recognize and handle these issues is also crucial for evaluation. Lastly, evaluation costs are high  Qin et al. (2023); Guo et al. (2024), as many studies rely on advanced large models for evaluation.
To address these shortcomings, we propose ACEBench, a comprehensive tool-use benchmark that includes the following categories:
Normal. Consists of fixed question-answer pairs and encompasses a variety of scenarios, including single-turn dialogues, multi-turn dialogues, and personalized scenario data. It also includes evaluations of atomic-level capabilities.
Special. Includes imperfect instructions, such as instructions containing incomplete parameters, incorrectly formatted parameters, or questions that are irrelevant to the capabilities of the candidate functions.
Agent. Encompasses real-world scenarios, abstracted to construct multi-turn, multi-step tool invocation scenarios, divided into multi-turn and multi-step cases depending on whether the user participates in the dialogue process.
The three categories above cover most of the tool usage scenarios for LLMs, and detailed explanations of each category can be found in Appendix A. Our main contributions are as follows:
  • Comprehensive Benchmark Evaluation. We propose a comprehensive benchmark for evaluating LLMs’ tool usage, covering various scenarios, including more fine-grained evaluation perspectives and assessments under imperfect instructions, and providing more stable evaluation metrics.
  • Sandbox Environment and Automated Evaluation System. We develop a sandbox environment construction scheme for multi-turn, multi-step tool invocation based on real-world scenario abstraction and build an end-to-end automated evaluation system.
  • Extensive Experimental Validation. Through extensive experiments, we demonstrate our benchmark provides a more comprehensive analysis with greater distinction, offering a clearer evaluation of LLMs’ tool usage.
Figure 1: Evaluation Dataset Construction Pipeline: API Synthesis Module (left), Dialogue Generation Module (middle), Quality Inspection Module (right).
2 Related Works
The emerging trend of leveraging LLMs’ tool-use capabilities in real-world applications underscores the need for comprehensive evaluations of their performance and effectiveness. Despite recent advancements, existing benchmarks for evaluating the tool-use capabilities of LLMs still have significant limitations
Stable ToolBench Guo et al. (2024) addresses the issue of unstable external APIs by employing a virtual API server, but its dependence on large models for evaluation results in high costs and scalability challenges. BFCL Yan et al. (2024) introduces a benchmark for tool use in multi-turn dialogue scenarios. Yet, it assembles dialogues from fixed content, failing to capture real-world interactions’ dynamic and adaptive nature. Similarly, Taubench Alanko et al. (2022) evaluates language agents’ ability to engage with human users while adhering to domain-specific rules. Still, its narrow focus on just two scenarios limits its generalizability across diverse tasks. HammerBench Wang et al. (2024a) improves upon this by incorporating datasets derived from popular mobile applications and merging dialogues to simulate typical question-answer trajectories. However, like BFCL, its multi-turn dialogues are simplistic concatenations of pre-defined content, which do not reflect the complexities of real-world conversational dynamics. In addition, some benchmarks Qin et al. (2023); Guo et al. (2024) rely on large language models (LLMs) for result evaluation, leading to high costs and unstable operations.
In contrast, our work addresses these limitations by expanding the scope of evaluation to encompass a broader range of tool usage scenarios. We propose a framework that simulates realistic multi-turn dialogue processes and enables end-to-end automated assessment, thereby reducing evaluation costs and improving scalability. A comparative analysis of ACEBench against recent benchmarks, as shown in Table 1, demonstrates its effectiveness in overcoming these challenges.
3 ACEBench
3.1 Dataset
We construct two versions of the dataset, one in Chinese and the other in English, ensuring an equal distribution of data types across both versions.
3.1.1 Data Generation
The data construction process is divided into two categories: (1) the creation of Agent data, as detailed in Appendix B.1, and (2) the creation of other types of data, which includes two primary steps: API synthesis and dialogue generation.
API Synthesis. We use real APIs from various real-world scenarios as reference during synthesis to enhance authenticity. To ensure the stability of the data, we use synthetic APIs to construct the evaluation dataset, referencing real-world APIs as a guide. We employ a self-evolution approach by building a hierarchical API context tree to ensure the generated APIs cover a wide range of domains and functionalities Liu et al. (2024b). Initially, we extract relevant information from technical documents to guide the API generation. As the process progresses, the context tree is gradually expanded, ultimately ensuring the depth and breadth of the generated APIs. The left part in Figure 1 illustrates the generation of APIs.
Dialogue Construction. As shown in the middle part in Figure 1, we use two different dialogue generation pipelines built on the constructed API pool from which three to six candidate APIs are selected for each evaluation instance. For most cases, APIs are chosen randomly. However, for instances requiring specific functionality (e.g., similar APIs or multi-turn scenarios), advanced methods, including graph-based sampling Wang et al. (2024d), are used. Simple cases or those with predefined functionality use a template-based generation, where a single generator produces dialogues to ensure consistency. We employ a multi-agent dialogue pipeline for more complex scenarios, where three agents (user, assistant, and tool) role-play to simulate real-world interactions. Both pipelines are supported by carefully hand-crafted examples to ensure comprehensive coverage and diversity. A detailed description for special data construction is provided in Appendix B.2.
3.1.2 Multi-Stage Data Verification
To address issues like mismatched answers or ambiguous criteria, we have implemented a multi-stage verification process, shown on the right part of Figure 1.
Automated Quality Inspection. The data first undergoes a rule-based quality inspection module, which evaluates four dimensions: clarity of API definitions, executability of function calls, accuracy of dialogues, and consistency of data samples, effectively filtering out formatting and spelling errors. Next, the data enters the model-based quality verification module, which uses LLMs to detect semantic errors, employing a voting mechanism to ensure consistency in evaluation.
Human Quality Inspection. In the initial evaluation, the dataset remaining after automated quality inspection is assessed by three LLMs to assist human experts in data screening. Valid data is retained, while potentially problematic data is placed in the error candidate pool (approximately 20% of the dataset). These flagged entries undergo a two-step expert review process, where two experts independently assess and suggest modifications, and a third expert consolidates feedback, revising problem statements, API definitions, and answers. The revised data is re-evaluated and manually verified, and three rounds of optimization are performed to ensure a high-quality dataset.
3.1.3 Data Analysis
To demonstrate the breadth and comprehensiveness of ACEBench, we provide a detailed analysis of its test case distributions. This includes an examination of API domains, the number of samples in each category, and the diversity of API patterns.
Figure 2: Distribution of APIs in terms of domains (Top 3 Subcategories for Each Category).
Domain of APIs. The ACEBench API boasts a comprehensive coverage of 8 major domains and 68 sub-domains, spanning various aspects of daily life, including technology, finance, entertainment, society, health, culture, environment, and others. It offers a rich collection of 4,538 APIs in both Chinese and English. The distribution of these APIs is visualized in the accompanying Figure 2.
Figure 3: Visualization of the data composition of ACEBench.
Data Composition. ACEBench consists of three categories of test samples: Normal, Agent, and Special, where each category is divided into several subcategories. The data composition is visualized in Figure 3, demonstrating a comprehensive coverage of tool-use capabilities, from simple single-turn tool invocations to complex multi-turn interactions involving users and environments. They include scenarios requiring multiple steps and interactions with the environment, as well as cases where tool calls are infeasible.
Figure 4: Distribution of Dialogue Turns and API Argument Numbers.
Number of turns and arguments. The test data in ACEBench covers a wide range of complexities. Specifically, we statistically analyzed the number of dialogue turns and the number of arguments in the called apis, which are visualized in Figure 4. The results show that the number of dialogue turns ranges from 1 to 8, encompassing most real-world scenarios. These samples with varying numbers of turns and arguments further form a test suite that covers a broader range of difficulties, allowing ACEBench to conduct a more granular evaluation of different models.
Figure 5: Overview of Evaluation Process: The left represents ’Normal’ evaluation, which involves AST parsing for function and parameter checks. The middle illustrates ’Special’ evaluation, focused on error type matching. The right shows ’Agent’ evaluation, which assesses final state and intermediate results through user-model interaction.
Table 2: Comprehensive Evaluation of Different Models on ACEBench for Chinese and English Combined (%).
Model Normal Special Agent Overall
Atom Single Turn Multi Turn Similar API Preference Summary
Closed-Source Large Language Models
GPT-4o 93.4 84.5 77.0 85.0 83.0 87.6 93.0 63.8 85.4
GPT-4-Turbo 93.2 84.8 77.5 86.0 86.0 88.0 86.7 67.5 84.5
Qwen-Max 91.2 80.5 68.0 83.0 83.0 84.2 74.0 64.3 78.4
GPT-4o-Mini 86.5 76.0 66.5 77.0 78.0 79.9 79.0 33.3 72.5
Gemini-1.5-Pro 84.5 76.8 64.5 80.0 78.0 79.0 78.7 25.5 70.7
Claude-3-5-Sonnet 76.9 72.5 62.5 71.0 72.0 72.9 77.4 39.5 68.9
Doubao-Pro-32k 79.8 55.5 58.0 76.0 66.0 70.7 55.0 25.0 59.4
Open-Source Large Language Models
Qwen2.5-Coder-32B-Instruct 90.2 81.0 71.0 83.0 81.0 84.1 80.7 60.8 79.6
DeepSeek-V3 91.5 84.0 77.0 83.0 83.0 86.5 73.0 34.5 74.8
Qwen2.5-72B-Instruct 86.8 80.3 69.5 83.0 81.0 82.1 75.7 45.0 74.7
Llama-3.1-70B-Instruct 82.5 68.3 63.5 79.0 68.0 75.5 38.3 42.3 60.4
Qwen2.5-7B-Instruct 76.0 60.3 58.5 72.0 67.0 69.4 47.0 13.8 54.8
DeepSeek-Coder-V2-Lite-Instruct 75.2 57.8 46.5 72.0 65.0 66.4 40.3 2.0 49.5
Qwen2.5-Coder-7B-Instruct 76.0 63.8 57.5 74.0 68.0 70.1 22.3 15.5 48.9
Watt-Tool-8B 85.7 69.3 55.5 79.0 64.0 75.6 6.0 2.8 45.7
Hammer2.1-7B 73.7 57.5 40.0 62.0 55.0 62.8 14.7 16.8 42.9
Llama-3.1-8B-Instruct 51.9 39.8 28.0 66.0 46.0 46.6 21.0 5.3 33.4
Phi-3-Mini-128k-Instruct 57.2 39.3 23.0 58.0 32.0 46.5 18.7 0.8 32.0
xLAM-7B-r 43.5 22.0 19.0 61.0 0.0 33.7 2.7 8.8 21.6
Llama-3.2-3B-Instruct 38.7 15.3 9.0 42.0 32.0 29.6 9.4 0.0 19.6
Hammer2.1-3B 22.4 11.5 3.5 40.0 20.0 18.7 1.0 1.5 11.3
3.2 Eval
In this section, we introduce the evaluation process, where different evaluation methods are applied based on the various data types. The inference prompt for evaluation can be found in Appendix D.
3.2.1 Normal Evaluation
As shown in the left part of Figure 5, we evaluate Normal Data by comparing the model’s function call outputs with the ground truth. We evaluate using AST parsing, starting by matching the function name to the target for exact correspondence, followed by verifying the consistency of parameter types, and finally confirming the accuracy of the parameter values. For cases with multiple possible correct answers, we use a candidate answer pool, and a match with any one of the candidates is considered correct. The Normal Data is evaluated using Accuracy, where a score of 1 indicates a complete match and 0 indicates a mismatch.
3.2.2 Special Evaluation
As shown in the middle part of Figure 5, the core goal of evaluating Special Data is to determine whether the model can accurately identify and point out the issue. Specifically, when the instruction is missing required parameters, the model needs to identify and indicate the missing parameter; when there is an error in the parameter, the model needs to correctly identify the error parameter; and when the task is irrelevant to the functions, the model needs to recognize that the function cannot process the request. In any of these cases, if the model correctly identifies and addresses the issue, the Accuracy is 1; otherwise, the Accuracy is 0.
3.2.3 Agent Evaluation
As shown in the right part of Figure 5, we evaluate the Agent Data by assessing the model’s ability to use tools in real-world interactions with humans, focusing on both result and process evaluation. There are two evaluation metrics:
End-to-End Accuracy is evaluated by comparing the instance attributes of the corresponding class with the target. If all attributes match exactly, the Accuracy is 1; otherwise, the Accuracy is 0.
Process Accuracy is determined by the consistency between the actual function call process and the ideal process. It is expressed as $\frac{n}{m}$, where $m$ represents the ideal function call process, and $n$ represents the degree of match between the actual and ideal processes.
3.2.4 Overall Accuracy
The Overall Accuracy is computed as a weighted sum of the accuracies for the Normal, Special, and Agent data types, where the weights are determined by the square roots of their respective sample sizes. The details can be found in Appendix F.
4 Experiments
In this section, we present a comprehensive set of experiments designed to evaluate the performance of LLMs on ACEBench.
Experimental Setup. In our evaluation, we examine seven closed-source LLMs, including the GPT-4 series Achiam et al. (2023), Qwen-Max Yang et al. (2024), Gemini-1.5-Pro Team et al. (2024), Claude-3.5-Sonnet Anthropic (2024), and Doubao-Pro-32K ByteDance (2025). Additionally, a wide range of open-source language models are assessed, such as the Qwen2.5 series Yang et al. (2024), Llama3 series Dubey et al. (2024), Phi-3-Mini Abdin et al. (2024), Deepseek-V3Liu et al. (2024a), and DeepSeek-Coder-V2 Zhu et al. (2024). Furthermore, four tool-learning-enhanced models were evaluated: Hammer2.1-3B, Hammer2.1-7B Lin et al. (2024), xLAM-7B-r Liu et al. (2024c), and Watt-Tool-8B Watt-AI (2024).
4.1 Main results and analysis
The comprehensive experimental results for the Chinese and English datasets are presented in Table 2, with detailed results for each language provided in Appendix F.1. We can draw the following important conclusions:
Performance Variations Between Top Open-source and Closed-source Models. The overall best performance still comes from closed-source models like the GPT-4 series, but the performance gap between certain open-source models, such as Qwen2.5-Coder-32B-Instruct and Qwen2.5-72B-Instruct, and existing closed-source models is narrowing. And as shown in Figure 6, the performance of these top models is generally similar in basic ability tasks, with the GPT-4 series models leading in most tasks, demonstrating strong intelligence. However, closed-source models outperform open-source models due to more refined internal optimizations and potentially larger scales. A noticeable performance gap emerges in mid-to-high-level tasks, such as Special and Agent, as these tasks require higher-level abilities, such as handling abnormal situations and strong multi-turn interaction and reasoning capabilities.
Figure 6: Performance of Top-3 Closed/Open-source models.
Loss of Generalization in Fine-Tuned Models. As shown in Table 3, models fine-tuned on specific datasets, such as Watt-Tool-8B Watt-AI (2024), xLAM-7B-r Liu et al. (2024c) and Hammer2.1-7B Lin et al. (2024), exhibit a significant decline in performance on the Special dataset. This decline can primarily be attributed to the fact that while fine-tuning improves the model’s performance on specialized tasks, it leads to a loss of general instruction-following ability.
Table 3: The Accuracy of Different Models on Special Data (%).
Model Incomplete Error Irrelevant
Llama-3.1-8B-Instruct 29.0 20.0 14.0
Watt-Tool-8B 7.0 1.0 10.0
Qwen2.5-7B-Instruct 26.0 36.0 79.0
xLAM-7B-r 1.0 3.0 4.0
Llama-3.2-3B-Instruct 13.0 12.0 3.0
Hammer2.1-3B 0.0 3.0 0.0
Performance Limitations of Large Models in Complex Tasks. As shown in Table 4, most models exhibit an end accuracy of less than 50% on Agent data tasks. This phenomenon can be attributed to the fact that completing a task in a dynamic environment, simulating real-world multi-turn interactions, requires more than just the model’s ability to perform individual tool operations. It also demands that the model effectively integrate contextual information during tool usage and consider the interdependencies between tool calls, making the task more challenging.
Table 4: Performance Evaluation of Different Models on Agent Data: PA represents Process Accuracy, EA represents End-to-End Accuracy (%).
Model Multi Turn Multi Step
EA PA EA PA
GPT-4-Turbo 50.0 66.06 85.0 89.5
DeepSeek-V3 31.5 54.5 37.5 53.0
Claude-3-5-Sonnet 21.5 41.5 57.5 76.5
DouBao-Pro-32k 20.0 45.5 30.0 47.5
Qwen2.5-7B-Instruct 15.0 28.0 12.5 15.5
Hammer2.1-7B 8.5 33.5 25.0 42.5
4.2 Error Analysis
Error Analysis of Normal Data. As shown in Figure 7, we observe from the error type distribution on Normal data that param value error dominate across all models. This highlights the models’ difficulty in generating specific values, likely due to limited contextual understanding and the complexity of numerical distributions. Output format error is the second most common, suggesting room for improvement in generating code that follows predefined formats and syntactic rules. These issues may stem from inconsistencies in training data and the models’ limited ability to learn rule-based generation. In contrast, function name and param type errors are less frequent, indicating that the models excel in matching function calls and handling data types. While the models show strong function invocation abilities, further improvements are needed in numerical generation and format compliance. Specific error examples for Normal data can be found in Appendix E.1.
Figure 7: Error type distribution across different model series on Normal Data.
Error Analysis of Special Data. As shown in Table 5, we identified two main types of model errors: The first type is "Error Detection", which refers to the model’s complete failure to detect issues in the user’s instructions or its inability to identify problems according to the prompt’s formatting requirements. The second type is "Error Correction," where the model detects the problem but provides unclear feedback. For example, the model might indicate that there is an issue, but fails to specify which parameter values are incorrect or what critical information is missing. Results show that most errors in special-type scenarios are caused by "Error Detection", highlighting a critical gap in the model’s problem-detection capabilities. This suggests that the model needs to learn not only simple tool invocation but also how to identify corresponding issues under imperfect instructions. Specific error examples for Special data can be found in Appendix E.2.
Table 5: Error type distribution across different model series on Special Data.
Model Error Detection Error Correction
Watt-Tool-8B 188 4
Hammer2.1-7B 172 7
Phi-3-mini-128k-instruct 143 15
Qwen2.5-3B-Instruct 130 36
xLAM-7B-r 195 1
Llama-3.1-8B-Instruct 145 6
Hammer2.1-3B 197 0
Llama-3.2-3B-Instruct 166 9
Error Analysis of Agent Data. Our analysis identifies three primary causes of Agent errors. First, function call errors occur when the model fails to select the appropriate function or provide parameters that do not meet the required specifications, reflecting a lack of understanding of tool-use capabilities and parameter constraints. Second, rule violations arise when the model disregards predefined scene rules, skipping necessary steps or breaking key task logic, highlighting deficiencies in its comprehension and execution. Finally, information mismanagement results from the model’s inability to correctly record or process contextual information during multi-turn interactions, leading to outputs that diverge from expectations. As shown in Figure 20, we illustrate an error caused by missing information.
4.3 Further Analysis
Scaling Law. We evaluated the performance of Qwen2.5-Coder (7B, 14B, 32B) and Qwen2.5-Instruct (3B, 7B, 14B, 32B, 72B) on the ACEBench dataset. The experimental results shown in Figure 8 demonstrate that performance improves significantly across various tasks as the model size increases, with particularly strong results on high-complexity tasks. This suggests that increasing the parameter size enables the model to achieve stronger task-solving and generalization capabilities.
Impact of Code Training. Based on the experimental results in Figure 8, we find that the score of Qwen2.5-Coder-7B is lower than the non-coder version, while the scores of Qwen2.5-Coder-14B and Qwen2.5-Coder-32B are higher than those of the non-coder versions of the same size. This indicates that code tuning becomes more impactful in improving model performance as the model size increases. For larger models (such as 14B and 32B), code tuning significantly enhances tool usage, especially in tasks that follow predefined rules and steps. However, for smaller models (such as 7B), despite code tuning, the differences between models are minimal, suggesting that smaller models have limited capacity and cannot fully leverage the advantages of code tuning.
Figure 8: Scaling Law of LLMs on ACEBench.
5 Conclusion
This paper introduces ACEBench, a comprehensive tool-use benchmark designed to evaluate the tool-use capabilities of Large Language Models (LLMs). It addresses key limitations of existing evaluation benchmarks, such as the lack of multi-turn dialogue assessments in real-world scenarios, the absence of fine-grained evaluations for parameter-type function calls, and the high costs associated with using large models for evaluation. The experimental results indicate that models fine-tuned on specific tool-use datasets to struggle with generalization when faced with complex or imperfect instructions, and code capabilities enhance the tool-use performance of large models. Through extensive experiments, we demonstrate the effectiveness of ACEBench in providing deeper insights into the tool-use abilities of various models.
Limitations
We acknowledge several limitations in our evaluation of ACEBench for assessing the tool-use capabilities of large language models. Firstly, while our test data is generated by large language models and various measures have been taken to ensure its authenticity and diversity, a gap remains when compared to data from real-world applications. This discrepancy may impact the evaluation of the model’s performance in real-world scenarios. Secondly, for the Agent data, the design of evaluation scenarios relies on manual construction, which somewhat limits the diversity and coverage of the evaluation framework.
Appendix A Detailed Descriptions of Test Cases
A.1 Data Categories Description
We divide the benchmark into three main categories: Normal, Special, and Agent. Below is a detailed description of each category.
Normal Data
The Normal Data consists of fixed question-answer pairs, where each question corresponds to a correct function call. It is categorized into the following categories: Single Turn, Multi Turn, Similar API, Preference and Atom. Data examples can be found in Appendix C.
Single Turn: There is only one-round interaction between the user and the assistant, and based on the number of function calls in the response, it is divided into single turn single function call and single turn parallel function call.
Multi Turn: There are multiple interactions between the user and the assistant. The conversation can be categorized into two types: (switch) The conversation progresses by changing topics. (adjust) The conversation evolves by refining or modifying the original question.
Similar APIs: The candidate APIs exhibit significant similarity, particularly in their focus on the same topic. This similarity presents a challenge for the assistant, requiring it to effectively distinguish between the APIs and accurately select the most appropriate ones.
Preference: In addition to the candidate APIs, the assistant is provided with supplementary user profile data. This type of information necessitates the assistant’s ability to mine user-specific factors, such as past interactions, interests, or other personalized attributes, to generate argument values.
Atom: Atom Data refers to a set of APIs that contain only specific parameter types, such as candidate functions where the parameters exclusively involve numbers, lists, etc. This design is intended to explore whether the type of function parameters affects the model’s ability to handle data filling. We have divided the Atom data into five types: number, enum, list, bool, and object.
Special Data
The Special Data refers to situations where the model is unable to resolve the problem posed in the instruction using the candidate functionsWang et al. (2024c). It is categorized into the following categories: Lose param, Error param, and Exceed Function Ability. Data examples can be found in Appendix C.2.
Incomplete: Refers to situations where the key information required for the function call is missing in the query, such as the absence of "required" parameters.
Error: Refers to situations where the instruction contains parameters or names that do not meet the required format or constraints, such as matching a specific pattern or being selected from a predefined list, causing the function call to fail.
Irrelevant: Refers to situations where the instruction exceeds the function’s capabilities, meaning none of the candidate functions can resolve the issue.
Agent Data
Agent Data refers to situations where completing a task in an environment built based on real-world scenarios typically requires multiple steps. In such cases, an advanced language model simulates the user role, replicating interactions found in real-world scenarios. This approach aims to evaluate the model’s ability to perform in complex interactions. We define a "multi-step" scenario as one in which the user appears only once in the entire conversation, while a "multi-turn" scenario is one where the user appears multiple times throughout the conversation. Data examples can be found in Figure C.3.
Appendix B The Construction of Data
B.1 Agent Data
As shown in Figure 9, the construction of Agent data can be summarized in the following steps:
First, through an in-depth analysis of real-world scenarios, extract key task requirements and modularize them into sub-scenarios in different domains (such as flight booking, food delivery platforms, and financial services), clearly defining the specific functional objectives of each module.
Second, implement logical abstraction for each function through code, designing core processes such as user authentication, cost calculation, payment processing, and information recording. This ensures the code logic’s scalability and robustness while comprehensively addressing exception handling.
Third, design interaction rules based on specific scenario requirements to standardize the interaction process between users and the Agent, such as verifying account, merchant information, and balance status in food delivery orders, and providing user guidance in exceptional cases.
Finally, combine real-world demands to design question formats and solutions, enabling the model to accurately meet user task requirements and achieve the desired outcomes.
Figure 9: The Construction of Agent Data. The left shows scenario sampling from real-world cases, the middle demonstrates the implementation of code tailored to specific scenarios, and the right presents examples of dialogue rules and question design for the scenarios.
B.2 Special Data
Incomplete
The special data which has incomplete instructions mainly refers to situations where key information is missing from the user’s instructions, causing the function to be called incorrectly. Our main prompt for constructing incomplete data is shown in  10. And we can extract the corresponding data and reference answers (missing data) from the generated dialogue.
{mdframed}
Please refer to the example and continue the dialogue based on the given tool definition and the beginning of the conversation. The requirements are as follows:
1 The user’s request is meaningful, requiring the use of one tool, and the tool will be called once.
2 Calling the tool to fulfill the user’s request still lacks one or more required parameters .
3 The assistant’s first response should be formatted as: [Thought] Missing parameters: Key1|Key2|…[Response] (follow-up question)". Replace Key1, Key2, etc., with specific parameter names, and replace the follow-up question with natural language content.
Here is an example for reference
<tool_definition>
"name": "calculate_triangle_area",
"description": "Given the base and height of a triangle, calculate its area.",
"required": ["base", "height"]
<dialogue_example>
[User]: I want to calculate the area of a triangle with a base of 5 and a height of 10.
[Assistant]: [Thought]  Missing parameters: base|height
[Response]:  Please provide both the base and the height of the triangle.
[User]: The base is 5, and the height is 10.
[Assistant]: <tool_usage>calculate_triangle_area|{"base": 5, "height": 10}</tool_usage>
Here is the dialogue continuation you need to write:
<Tool Definition>
{tool_definition}
Figure 10: Example of Incomplete Data construction.
Next, we can convert the conversation we obtained into data. Here is a specific example.
{mdframed}
candidate function
  "name": "book_flight",
  "description": "Flight booking tool...",
  "arguments": {},
  "required": ["from_city_name", "to_city_name", "depart_date"]
obtained dialouge
<user> I plan to travel to Beijing. Help me book a flight to Beijing on June 1st.
</user> [Thought] Missing parameters: from_city_name
[Response] May I know your departure city you want to book?
<user> Oh, Im in Shenzhen, economy class X
converted data
[question] I plan to travel to Beijing. Help me book a flight to Beijing on June 1st.
[answer] Miss necessary parameter (from_city_name) from the api (book_flight)
Error
The special data which has error instructions mainly refers to situations where the instruction contains parameters or names that do not meet the required format or constraints
{mdframed}
Please refer to the example and continue the dialogue based on the given tool definition and the beginning of the conversation. The requirements are as follows:
1 The user’s request is specific and meaningful, requiring the use of one tool, and the tool should be called once.
2 The user’s first question contains incorrect parameters (i.e., it does not conform to the function definition’s pattern or format requirements).
3 The assistant’s first reply should accurately identify the error and ask the user for more precise parameters, formatted as: ’Thought Error in parameter: Key1 Response (Follow-up question content)’. Replace ’Key’ with the specific parameter name and provide a natural, real-language follow-up question."
Here is an example for reference
<tool_definition>
{"name": "TennisTeamDetails",
  "description": "Retrieve detailed information about a tennis team by its name.",
  "parameters": {
    "properties": {
      "team_name": {......
[…]
<dialogue_example>
<user> Can you retrieve the details of the tennis team named ’Team@1234’?
</user> [Thought] Error Parameter: ’Team@1234’
[Response]  The team name you provided does not meet the naming criteria.
<user> Oh, I made a mistake. The team name is TeamABC.
</user> <tool_usage>TennisTeamDetails|{"team_name": "TeamABC"}</tool_usage>"
Here is the dialogue continuation you need to write:
<Tool Definition>
{tool_definition}
Figure 11: Example of Incomplete Data construction.
Next, we can convert the conversation we obtained into data. Here is a specific example.
{mdframed}
candidate function
{ "name": "FootballTeamDetails",
  "description": "Retrieve detailed information about a football team by its name.",
  "parameters": {
      "team_name": {......
[…]
obtained dialouge
<user> I want to know information about the football team football\$156.
</user> [Thought] Error Parameter: football\$156.
[Response] The team name you provided does not meet the naming criteria.
<user> ......
converted data
[question] I want to know information about the football team footbaoo\$156.
[answer] There is incorrect value (football\$156) for the parameters (team_name).
Irrelevant
The construction method for Irrelevant data is similar to that of Normal data, where we simply remove the correct API from the candidate APIs of the data.
Appendix C Examples of Dataset
C.1 Normal Examples
{mdframed}
Please assess the system with asset ID 12345, vulnerability ID 67890, with a severity level of high, and a potential threat that may lead to data leakage.
(candidate function)
  "name": "Threat_Vulnerability_Assessment_Tool",
  "description": "Assess potential security threats and provide mitigation suggestions",
  "parameters": ......
  "name": "userPrivacySettingsManager",
  "description": "Manage user’s privacy settings",
  "parameters": ......
Figure 12: Example of Normal Single Turn Data.
{mdframed}
User: I want to analyze the pest and disease risks faced by my rice crop, which is currently in the tillering stage.
system: Please provide the current temperature and humidity conditions.user: The temperature is 28°C, and the humidity is 80%.
tool:
[PestInfestationRiskAnalyzer(
    crop_information={"crop_type": "rice", "growth_stage": "tillering"},
    climate_conditions={"temperature": 28, "humidity": 80})]
system: Based on the analysis, your rice is facing a moderate risk of pests and diseases under the current climatic conditions. Please take necessary preventive measures.
User: I also want to know the maintenance requirements for my production equipment. The device ID is Device123.
system: Please provide the sensor readings of the device, including temperature, vibration, and humidity.
User: The device currently reads a temperature of 75°C, vibration level of 5, and surrounding humidity of 50%.
(candidate function)
  "name": "PestInfestationRiskAnalyzer",
  "description": "Analyze the pest and disease risk faced by the farmland ",
  "parameters": ......
  "name": "Predictive_Maintenance_AI_Tool",
  "description": "Predicting the maintenance needs of equipment using sensor data",
  "parameters": ......
  "name": "PersonalityTraitAnalyzer",
  "description": "Analyze an individual’s personality traits",
  "parameters": ......
Figure 13: Example of Normal Multi Turn Data.
{mdframed}
Can you update my preferred notification method and also check if my current email address and home location are properly updated in the system?
(candidate function)
  "name": "updateOrderStatusAlerts",
  "description": "Sends automated alerts to users regarding the status of their current orders.",
  "parameters": ......
  "name": "submitProductReview",
  "description": "Allows users to submit a review for a product they have purchased.",
  "parameters": ......
  "name": "updateUserProfile",
  "description": "Updates the user’s profile information based on provided data.",
  "parameters": ......
  "basic_features": {
    "UserName": "Michael Smith",
    "UserEmail": "mike.smith@example.com",
    "UserHomeLocation": "Los Angeles, CA",
    "UserBirthday": "1978-04-23",
    "UserLanguage": "Spanish",
    "UserTimeZone": "PST",
  "user_history": {
    "shopping": [
      "Searched for ’Nike running shoes’ on app",
      "Added Nike Air Max to cart",
      "Checked coupon availability for Nike products",
      "Filtered search by ’Outdoor Equipment’ category",
      "Selected ’High spending’ filter for items over \$500",
    "takeout": [
      "Ordered Chicken Fajitas on the takeout app for lunch",
      "Opted to receive promotional deals via phone calls",
      "Chose Debit Card ending in 5678 for payment on the takeout app",
Figure 14: Example of Normal Preference Data.
{mdframed}
My baby has had a visible vein on her nose for 5 days, and she’s been crying a lot with a decreased appetite. Can you help?
(candidate function)
  "name": "baby_health_check_A",
  "description": "Checks the common reasons for baby’s persistent vein visibility on the nose
      and suggests actions. This API considers factors like skin thinness, crying, or overexertion",
  "parameters": ......
  "name": "baby_health_check_B",
  "description": "Examines baby’s vein visibility and provides a  recommendation to see a doctor.
      Focuses on persistent visibility and associated symptoms",
  "parameters": ......
Figure 15: Example of Similar API Data.
{mdframed}
I need a design for my new website. It’s for a technology company focusing on user engagement.
(candidate function)
  "name": "WebDesignAssistant_generateDesign",
  "description": "Generates a website design based on industry and user experience focus.",
  "parameters": {
    "type": "object",
    "properties": {
      "industry": {
        "description": "The industry for which the website is being designed.",
        "type": "string",
        "enum": [
          "Technology",
          "Healthcare",
          "Education",
          "Finance"
        "default": "Technology"
      "userExperience": {......}
    "required": ["industry", "userExperience"]
Figure 16: Example of Atom (enum) Data.
C.2 Special Example
{mdframed}
I’m considering relocating my business to the Middle East. Can you provide me with a list of major cities?
(candidate function)
{"name": "Get_Middle_East_Cities",
  "description": "Retrieves a list of cities in the Middle East, sorted by overall score by default.",
  "parameters": {
    "properties": {
      "sort": {
        "description": "The sorting order for the list of cities.",
        "type": "string",
        "enum": ["asc", "desc"],
    "required": ["sort"]
Figure 17: Example of Special Lose Param Data.
{mdframed}
Can you retrieve the tennis team details for the team named ’Team@1234’?”
(candidate function)
{"name": "TennisTeamDetails",
  "description": "Retrieve detailed information about a tennis team by its name.",
  "parameters": {
    "type": "object",
    "properties": {
      "team_name": {
[…]
    "required": ["team_name"]
Figure 18: Example of Special Error Param Data.
{mdframed}
Could you help me find available restaurants in New York City?
(candidate function)
  "name": "Get_Weather_Report",
  "description": "Retrieve the current weather report for a specified location",
  "parameters": ......
  "name": "GetTravelDestinationInfo",
  "description": "Retrieves information about a specific travel destination",
  "parameters": ......
Figure 19: Example of Special Exceed Function Ability Data.
C.3 Agent Example
Figure 20: Example of Agent Data.
Figure 21: Example of Agent Data.
Appendix D Evaluation Inference Prompts
D.1 Normal Prompt
{mdframed}
You are an AI assistant with the role name "assistant." Based on the provided API specifications and conversation history from steps 1 to t, generate the API requests that the assistant should call in step t+1. The API requests should be output in the format [ApiName(key1=’value1’, key2=’value2’, …)], replacing ApiName with the actual API name, key1, key2, etc., with the actual parameter names, and value1, value2, etc., with the actual parameter values. The output should start with a square bracket "[" and end with a square bracket "]". If there are multiple API requests, separate them with commas, for example:[ApiName(key1=’value1’,key2=’value2’,…),ApiName(key1=’value1’,key2=’value2’, …), …]. Do not include any other explanations, prompts, or API call results in the output. If the API parameter description does not specify otherwise, the parameter is optional (parameters mentioned in the user input need to be included in the output; if not mentioned, they do not need to be included). If the API parameter description does not specify the required format for the value, use the user’s original text for the parameter value. If the API requires no parameters, output the API request directly in the format [ApiName()], and do not invent any nonexistent parameter names.
{time}
Role Descriptions:
user: User
assistant: The AI assistant role that makes API requests
tool: Provides the results returned from tool calls
API Specifications:
{function}
Figure 22: System Prompt for Nomal (Except Prefernce).
{mdframed}
You are an AI assistant, and your role is called assistant. Based on the given API description, dialogue history 1..t, and character profile, generate the API requests that the assistant should call in step t+1. The API requests should be output in the format [ApiName(key1=’value1’, key2=’value2’, …)], where ApiName is replaced with the actual API name, and key1, key2, etc., are replaced with the actual parameter names, and value1, value2 are replaced with the actual parameter values. The output should start with a "[" and end with a "]". If there are multiple API requests, they should be separated by commas, e.g., [ApiName(key1=’value1’, key2=’value2’, …), ApiName(key1=’value1’, key2=’value2’, …), …]. Do not output any other explanations, hints, or results of the API calls in the output. If the API parameter description does not specify special instructions, the parameter is optional (parameters mentioned in the user input or character profile should be included in the output, and if not mentioned, they should not be included). If the API parameter description does not specify the format for the parameter value, the parameter value should be taken from the user’s original text or character profile. If the API requires no parameters, the API request should be output as [ApiName()], with no fabricated parameter names.
Character Profile:
{profile}
Role Descriptions:
user: User
assistant: The AI assistant role that makes API requests
tool: Provides the results returned from tool calls
API Specifications:
{function}
Figure 23: System Prompt for Nomal (Prefernce).
D.2 Special Prompt
{mdframed}
You are an AI system with the role name "system." Based on the provided API specifications and conversation history from steps 1 to t, generate the API requests that the system should call in step t+1. Below are two specific scenarios:
1. When the information provided by the user is clear and unambiguous, and the problem can be resolved using the list of candidate functions:
- If the API parameter description does not specify the required format for the value, use the user’s original text for the parameter value.
- API requests should be output in the format [ApiName(key1=’value1’, key2=’value2’, …), ApiName(key1=’value1’, key2=’value2’, …), …], replacing ApiName with the actual API name, key1, key2, etc., with the actual parameter names, and value1, value2, etc., with the actual parameter values. The output should start with a square bracket "[" and end with a square bracket "]". At this time, the output must not contain any other content.
2. When the information provided by the user is unclear, incomplete, or incorrect, or the user’s question exceeds the capabilities of the provided functions, you need to clearly point out these issues. The following is your strategy:
(1) If the user’s instructions include the key details required to call the API, but the type or form of the parameter values does not match the API’s definitions, ask in-depth questions to clarify and correct the details. The output format should be: ["There is incorrect value (value) for the parameters (key) in the conversation history."]
(2) If the user’s instructions lack the key details required by the API, ask questions to obtain the necessary information. The output format should be: ["Missing necessary parameters (key1, key2, …) for the api (ApiName)"], replacing key1, key2 with the names of the missing parameters and ApiName with the actual API name.
(3) If the user’s request exceeds the current capabilities of your APIs, inform them that you cannot fulfill the request. The output format should be: ["Due to the limitations of the function, I cannot solve this problem."]
Note: The above steps have a priority order. You need to first determine whether scenario (1) applies. If it does, output according to the requirements in (1). Pay attention to distinguishing between scenarios (1) and (2).
{time}
Role Descriptions:
user: User
assistant: The AI assistant role that makes API requests
tool: Provides the results returned from tool calls
API Specifications:
{function}
Figure 24: System Prompt for Special.
D.3 Agent Prompt
{mdframed}
The current time is June 11, 2024, 16:00 (Beijing Time). As a simulated mobile assistant agent, you can help users send text messages, add reminders, and order takeout.
Text messages
Sending Text Messages
(1)Before sending a text message, the agent must first obtain the sender and recipient of the message.(2)When the memory is full and needs to delete messages, you need to ask the user: "Memory is full, which message would you like to delete?"
Viewing Text Messages (1)Before viewing text messages, the agent must first log into the device via login_device().(2)Before viewing text messages, the agent must first obtain the sender and recipient of the messages.(3)After viewing text messages, the agent needs to ask the user if they want to add the message content to a reminder.(4)After viewing text messages, the agent needs to ask the user if they want to reply to the message.(5)If the message content involves takeout, the agent needs to ask if the user wants to order takeout based on the message content.
Reminders
Adding Reminders
(1)Before adding a reminder, you should obtain the content and title of the reminder. The reminder time defaults to the current time.(2)If the reminder to be added is the content of a specific message, the agent needs to first view the message content.
Viewing Specific Reminders by Title:After viewing a specific reminder by title, you need to ask the user if they want to complete the tasks within it.
Order takeout
Ordering Takeout
(1)Before ordering takeout, the agent needs to obtain the user’s takeout platform account and password, and log in using login_food_platform().(2)If the merchant, product, and quantity for the order are not initially provided, you need to ask the user.(3)When encountering takeout from different merchants, you need to order them one by one.(4)If the balance is insufficient, you need to inform the user "Insufficient balance" and ask if they want to change the order.
Function Calls
When a function call is needed, please strictly adhere to the above format requirements:
(1)[ApiName(key1=’value1’, key2=’value2’, …)], Please remember that the function call must start with [ and end with ]
(2)You need to promptly feedback the task execution status to the user and do not repeatedly call the same function. When you believe the current task is completed, respond with "finish conversation" to end the dialogue.
Figure 25: System Prompt for Agent.
Appendix E Error Examples
E.1 Error Examples of Normal Data
{mdframed}
I want to understand the symmetry in Escher’s Ẅaterfallärtwork. Please provide a detailed symmetry analysis.
{"Escher_Artwork_Analysis_Tool": {
    "artwork_analysis": [
      {"artwork_id": "Waterfall",
       "analysis_type": "Symmetry Analysis"}]}}
(model output)
{“FontAdvisor_recommendDisplayFonts”:{……}}
Figure 26: Error Example of Noraml Wrong Function.
{mdframed}
Can you check if our current automation system with system ID 2435 is compatible with the new device model XR10?
{"AutomationSystemCompatibilityCheck":{ "systemID":2435,"newDeviceModel":"XR10"}}
(model output)
{“AutomationSystemCompatibilityCheck”:{ “systemID”:"2435", "newDeviceModel":"XR10"}}
Figure 27: Error Example of Noraml Param Type.
{mdframed}
I would like to assess the GPS accuracy of my vehicle. The device ID is GPS-7524. Sorry, I made a mistake earlier—it should be GPS-7526.
{ "SignalAccuracyEvaluator":{ "device_id":"GPS-7526“ }}
(model output)
{"SignalAccuracyEvaluator":{ "device_id":"GPS-7524“ }}
Figure 28: Error Example of Noraml Param Value.
{mdframed}
I am currently working on a research paper about the state of forests, and I need an assessment of the vegetation diversity in the Amazon rainforest. Please provide me with a detailed report.
{"EcoDiversityMonitor_assessVegetation": {
    "area_identifier": "Amazon rainforest",
    "detailed_report": true}}
(model output)
[user,area_identifier=’Amazonrainforest’,detailed_report=True]
Figure 29: Error Example of Wrong Output Format.
E.2 Error Examples of Special Data
{mdframed}
I‘m considering relocating my business to the Middle East. Can you provide me with a detailed list of major cities there, sorted by internet speed?
Missing necessary parameters (sort) for the api (Get_Middle_East_Cities)
(model output)
  "Get_Middle_East_Cities": {
    "sort": "desc",
    "sort_by": "internet_speed",
    "size": 10,
    "page": 1
(The question did not mention the sort method)
Figure 30: Error Example of Special Error Blindness.
{mdframed}
I‘m considering relocating my business to the Middle East. Can you provide me with a detailed list of major cities there, sorted by internet speed?
Missing necessary parameters (sort) for the api (Get_Middle_East_Cities)
(model output)
Missing necessary parameters (sort by) for the api (Get_Middle_East_Cities)
(Find missing necessary parameters, but not correctly identified)
Figure 31: Error Example of Special Vague Error Reporting.
Appendix F Formula for Overall Accuracy
The formula for calculating the Overall Accuracy can be expressed as:
All $\text{All Acc}=A\cdot\text{Acc}_{\text{Normal}}+B\cdot\text{Acc}_{\text{Special}}+C\cdot\text{Acc}_{\text{Agent}}$
where the coefficients $A$, $B$, and $C$ are defined as:
$A=\frac{\sqrt{n_{\text{Normal}}}}{\sqrt{n_{\text{Normal}}}+\sqrt{n_{\text{Special}}}+\sqrt{n_{\text{Agent}}}}$
$B=\frac{\sqrt{n_{\text{Special}}}}{\sqrt{n_{\text{Normal}}}+\sqrt{n_{\text{Special}}}+\sqrt{n_{\text{Agent}}}}$
$C=\frac{\sqrt{n_{\text{Agent}}}}{\sqrt{n_{\text{Normal}}}+\sqrt{n_{\text{Special}}}+\sqrt{n_{\text{Agent}}}}$
- where $n_{\text{Normal}},n_{\text{Special}},n_{\text{Agent}}$ are the sample sizes for the Normal, Special, and Agent data types, and Accuracy refers to the accuracy of each respective category.
F.1 Detailed Results
Table 6: Accuracy Evaluation of Different Models on English Data (%).
Model Normal Special Agent Overall
Atom Single Turn Multi Turn Similar API Preference Summary
Closed-Source Large Language Models
GPT-4o 90.0 78.0 68.0 80.0 78.0 82.5 92.7 56.0 81.1
GPT-4-Turbo 90.7 80.5 69.0 80.0 88.0 84.2 82.0 62.5 80.3
Qwen-Max 88.0 75.0 61.0 74.0 82.0 79.7 74.0 60.0 75.1
GPT-4o-Mini 84.3 73.5 59.0 74.0 72.0 76.4 76.7 27.5 68.9
Gemini-1.5-Pro 82.3 73.0 61.0 74.0 72.0 75.7 77.3 26.0 68.5
Claude-3-5-Sonnet 66.7 64.0 46.0 58.0 68.0 62.2 72.7 44.0 62.2
Doubao-Pro-32k 75.3 58.0 52.0 70.0 54.0 66.3 50.7 26.5 56.0
Open-Source Large Language Models
Qwen2.5-Coder-32B-Instruct 86.0 73.5 59.0 76.0 72.0 77.4 80.0 50.0 73.9
DeepSeek-V3 88.0 77.5 63.0 76.0 78.0 80.3 72.7 34.0 71.1
Qwen2.5-72B-Instruct 81.3 74.5 64.0 76.0 80.0 76.8 74.0 37.5 70.0
Llama-3.1-70B-Instruct 83.7 71.5 61.0 74.0 66.0 75.6 29.3 41.0 57.9
Qwen2.5-7B-Instruct 70.3 57.0 49.0 62.0 58.0 62.8 49.3 15.0 51.8
Qwen2.5-Coder-7B-Instruct 73.3 63.5 52.0 70.0 58.0 66.6 25.3 18.5 48.1
DeepSeek-Coder-V2-Lite-Instruct 71.7 58.0 50.0 62.0 60.0 64.0 39.3 2.5 47.9
Watt-Tool-8B 84.7 71.5 57.0 70.0 62.0 74.8 2.0 1.5 44.0
Hammer2.1-7B 71.3 62.5 43.0 64.0 52.0 62.9 3.3 15.0 39.6
Phi-3-Mini-128k-Instruct 66.3 49.0 31.0 58.0 32.0 54.0 12.0 0.0 34.4
MLlama-3.1-8B-Instruct 51.0 49.5 28.0 60.0 56.0 48.1 15.3 6.5 32.9
xLAM-7B-r 61.7 42.0 32.0 66.0 0.0 48.7 4.0 10.0 30.8
Llama-3.2-3B-Instruct 31.7 21.5 9.0 34.0 32.0 26.4 8.7 0.0 17.6
Hammer2.1-3B 32.7 14.0 7.0 36.0 32.0 25.5 0.7 1.5 15.2
Table 7: Accuracy Evaluation of Different Models on Chinese Data (%).
Model Normal Special Agent Overall
Atom Single Turn Multi Turn Similar API Preference Summary
Closed-Source Large Language Models
GPT-4o 96.7 91.0 86.0 90.0 88.0 92.7 93.3 71.5 89.6
GPT-4-Turbo 95.7 89.0 86.0 92.0 84.0 91.7 91.3 72.5 88.6
Qwen-Max 94.3 86.0 75.0 92.0 84.0 88.7 74.0 68.5 81.7
GPT-4o-Mini 88.7 78.5 74.0 80.0 84.0 83.4 81.3 39.0 76.0
Claude-3-5-Sonnet 87.0 81.0 79.0 84.0 76.0 83.5 82.0 35.0 75.6
Gemini-1.5-Pro 86.7 80.5 68.0 86.0 84.0 82.2 80.0 25.0 72.8
Doubao-Pro-32k 84.3 53.0 64.0 82.0 78.0 75.0 59.3 23.5 62.8
Open-Source Large Language Models
Qwen2.5-Coder-32B-Instruct 94.3 88.5 83.0 90.0 90.0 90.8 81.3 71.5 85.3
Qwen2.5-72B-Instruct 92.3 86.0 75.0 90.0 82.0 87.3 77.3 52.5 79.3
DeepSeek-V3 95.0 90.5 91.0 90.0 88.0 92.6 73.3 35.0 78.5
Llama-3.1-70B-Instruct 81.3 65.0 66.0 84.0 70.0 75.3 47.3 43.5 62.9
Qwen2.5-7B-Instruct 81.7 63.5 68.0 82.0 76.0 75.9 44.7 12.5 57.8
DeepSeek-Coder-V2-Lite-Instruct 78.7 57.5 43.0 82.0 70.0 68.8 41.3 1.5 51.1
Qwen2.5-Coder-7B-Instruct 78.7 64.0 63.0 78.0 78.0 73.5 19.3 12.5 49.6
Watt-Tool-8B 86.7 67.0 54.0 88.0 66.0 76.3 10.0 4.0 47.4
Hammer2.1-7B 76.0 62.5 37.0 60.0 58.0 62.7 26.0 18.5 46.1
Llama-3.1-8B-Instruct 52.7 30.0 28.0 72.0 36.0 45.0 26.7 4.0 33.8
Phi-3-Mini-128k-Instruct 48.0 29.5 15.0 58.0 32.0 38.9 25.3 1.5 29.5
Llama-3.2-3B-Instruct 45.7 9.0 9.0 50.0 32.0 32.7 10.0 0.0 21.6
xLAM-7B-r 25.3 2.0 6.0 56.0 0.0 18.7 1.3 7.5 12.3
Hammer2.1-3B 12.0 9.0 0.0 44.0 8.0 11.8 1.3 1.5 7.4